1. What type of demand curve does a perfectly competitive firm face? Why?

2. A perfectly competitive firm has the following fixed and variable costs in the short run. The market price for the firm’s product is $150.
       Output   FC   VC   TC   TR   Profit/Loss
               0   $100   $    0   ___   ___   ___
       1       100   100   ___   ___   ___
       2       100   180   ___   ___   ___
       3       100   300   ___   ___   ___
       4       100   440   ___   ___   ___
       5       100   600   ___   ___   ___
       6       100   780   ___   ___   ___
       a.   Complete the table.
       b.   At what output rate does the firm maximize profit or minimize loss?
       c.   What is the firm’s marginal revenue at each positive level of output? Its average revenue?
Answer #1

1. What type of demand curve does a perfectly competitive firm face? Why?

Answer:

perfectly elastic demand (Horizontal demand curve).

Why?

because they are price taker. so they charge market price. at any quantity price will be same. so it has perfectly Elastic demand curve that is horizontal demand curve.

2.

a.   Complete the table.

Q FC VC TC TR profit
0 100 100 0 -100
1 100 100 200 150 -50
2 100 180 280 300 20
3 100 300 400 450 50
4 100 440 540 600 60
5 100 600 700 750 50
6 100 780 880 900 20

TR=Price*Q

TC=TVC+TFC

profit=TR-TC

b.   At what output rate does the firm maximize profit or minimize loss?

Answer:

4 unit.

explanation: firm maximizes its profit where, MR=MC. but in this case MR is not equal to MC at any quantity so we will select the quantity where MR is greater than MC before MC starts becoming greater than MR. so here at quantity 4 MR is Greater than MC before MC starts becoming greater than MR.

Q FC VC TC TR profit MR MC
0 100 100 0 -100
1 100 100 200 150 -50 150 100
2 100 180 280 300 20 150 80
3 100 300 400 450 50 150 120
4 100 440 540 600 60 150 140
5 100 600 700 750 50 150 160
6 100 780 880 900 20 150 180

MC=change in TC/change in quantity

MR=change in TR/change in quantity

c.   What is the firm’s marginal revenue at each positive level of output? Its average revenue?

Q FC VC TC TR profit MR MC AR
0 100 100 0 -100
1 100 100 200 150 -50 150 100 150
2 100 180 280 300 20 150 80 150
3 100 300 400 450 50 150 120 150
4 100 440 540 600 60 150 140 150
5 100 600 700 750 50 150 160 150
6 100 780 880 900 20 150 180 150

MR=change in TR/change in quantity

AR=TR/Q

Note : in perfect competition AR=MR=price.